# Context-Free Recognition with Transformers

### v1

@@ Abstract
Transformers excel on tasks that process well-formed inputs according to some grammar, such as natural language and code. However, it remains unclear how they can process grammatical syntax. In fact, under standard complexity conjectures, standard transformers cannot recognize context-free languages (CFLs), a canonical formalism to describe syntax, or even regular languages, a subclass of CFLs (merrill-etal-2022-saturated). merrill-2025littledepthgoeslong show that ${{\mathcal{O}}}(\log({{n}}))$ looping layers (w.r.t. input length n{{n}}) allows transformers to recognize regular languages, but the question of context-free recognition remained open. In this work, we show that looped transformers with ${{\mathcal{O}}}(\log({{n}}))$ looping layers and ${{\mathcal{O}}}({{n}}^{6})$ padding tokens can recognize all CFLs. However, training and inference with ${{\mathcal{O}}}({{n}}^{6})$ padding tokens is potentially impractical. Fortunately, we show that, for natural subclasses such as unambiguous CFLs, the recognition problem on transformers becomes more tractable, requiring ${{\mathcal{O}}}({{n}}^{3})$ padding. We empirically validate our results and show that looping helps on a language that provably requires logarithmic depth. Overall, our results shed light on the intricacy of CFL recognition by transformers: While general recognition may require an intractable amount of padding, natural constraints such as unambiguity yield efficient recognition algorithms.
@@ 1 Introduction
Transformers are proficient at many natural language (qin2024largelanguagemodelsmeet) and coding (jiang2024surveylargelanguagemodels) tasks, both of which involve processing hierarchical structures. Classically, the ability to process hierarchically nested structures is closely connected to the ability to model context-free languages (CFLs). Analysis of internal representations—syntactic probing—has shown that transformers learn to encode syntactic features relevant for parsing, the task of extracting the syntactic structure of a sentence (hewitt-manning-2019-structural; arps2022probingconstituencystructureneural; zhao-etal-2023-transformers). However, it is unclear what classes of syntax transformers can provably represent, and how CFL recognition can be implemented internally. To this end, we study whether transformers can correctly determine the grammaticality of a sentence according to a context-free grammar.
@@
The problem of determining whether an input is grammatical can be stated as the recognition problem for context-free grammars (CFGs): Given a CFG $\mathcal{G}$}}, can a string $\bm{w}$}} be generated by $\mathcal{G}$}}? Several foundational serial parsing algorithms (earley-parsing; cocke; kasami1965AnER; younger) solve this problem. However, such serial procedures cannot be naturally implemented by transformers due to their highly parallel, fixed-depth structure. Even regular languages, a strict subset of CFLs, cannot be recognized by fixed-depth transformers under the standard complexity conjecture ${\textsc{TC}\textsuperscript{0}}\subsetneq{\textsc{NC}\textsuperscript{1}}$: Regular language recognition is complete for NC1 (barrington_monoids) while fixed-depth transformers fall in TC0 (merrill-etal-2022-saturated; chiang2025transformersuniformtc0). Looping layers help: $\log({{n}})$ looping layers (where n{{n}} is the input length) allow transformers to recognize regular languages (merrill-2025littledepthgoeslong). However, the question of whether logarithmic looping enables CFL recognition remains. In this work, we address it by analyzing the difficulty of recognizing various CFL classes by transformers. We conceptualize the difficulty in terms of extra resources needed: looping layers and appending blank padding tokens (merrill2025exactexpressivepowertransformers).
@@
While general CFL recognition cannot be implemented by fixed-depth transformers under standard complexity conjectures, our first result shows via a direct construction that it can be expressed by looping layers ${{\mathcal{O}}}(\log({{n}}))$ times and with ${{\mathcal{O}}}({{n}}^{6})$ padding tokens. To the best of our knowledge, this constitutes the first proof of general CFL recognition by transformers. We then ask whether simpler classes of CFLs can be recognized by transformers with fewer resources. We find that the answer is affirmative: We show that natural subclasses of CFLs can be recognized by simpler transformers. In particular, we identify unambiguity and linearity as two key properties that make CFL recognition more tractable. Unambiguous CFLs, characterized by strings having at most one possible parse, allow for recognition with reduced padding but more looping. This aligns with transformers’ struggles to parse ambiguous grammars in practice (khalighinejad-etal-2023-approximating). Furthermore, imposing linearity (where each grammar rule has at most one non-terminal on its right-hand side) reduces the amount of looping and padding required for recognizing unambiguous CFLs. We empirically test when looping helps generalization and find it to increase the performance on a CFL known to require ${{\mathcal{O}}}(\log({{n}}))$ time on parallel models of computation, namely the language of variable-free Boolean formulas (buss-bfvp).
@@
In summary, we leverage theory on parallel recognition of CFLs to show that logarithmically-looped transformers can recognize CFLs, characterizing the padding requirements for different relevant subclasses. These results imply that, in order to recognize CFLs, transformers require exponentially less depth than what would be needed to implement a serial parsing algorithm like CKY. While this comes with increased space (padding) requirements in the general case, we show the space can be reduced for natural CFL subclasses. Our results open the avenue for tighter upper bounds on the required space for CFL recognition.
@@ 2.1 Context-free grammars
The language of a CFG $\mathcal{G}$}} is the set ${\mathbb{L}}({{\mathcal{G}}})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\{{{\bm{w}}}\in{{{{\Sigma}}^{*}}}\mid{{{{\mathrm{S}}}}}{{\overset{*}{\Rightarrow}}}{{\bm{w}}}\}$.
@@ 2.2 Transformers
We consider the idealization of transformers from merrill-2025littledepthgoeslong; merrill2025exactexpressivepowertransformers. In short, [^1] We refer to App. A for more details on the transformer model. we study average hard attention transformers (AHATs), where the attention normalization function returns a uniform average of the values of tokens that maximize the attention score. The transformers use multi-pre-norm, where the layer normalization is applied before the residual connection on either the entire hidden state or on distinct subsets thereof (merrill2024expressivepowertransformerschain). We further assume logarithmic-precision arithmetic, where computations are performed with ${{\mathcal{O}}}(\log({{n}}))$ bits for an input of size n{{n}}. Coupling AHATs and log-precision unlocks useful gadgets such as storing string indices, counting symbol occurrences across the string and performing equality checks of values stored in residual streams at separate positions (merrill2024expressivepowertransformerschain). We assume input strings to the transformer are augmented with both a beginning-of-sequence (bos) and end-of-sequence (eos) token. Denote by ${{{\bm{x}}}}_{{{\textsc{eos}}}}^{L}$ the contextual representation of eos at end of the forward pass of the transformer. We apply a linear classifier to ${{{\bm{x}}}}_{{{\textsc{eos}}}}^{L}$ to determine string acceptance.
@@ 3 Recognizing general CFLs with transformers
Our goal is to recognize a CFL represented by a grammar in CNF (Def.˜2.4) with start symbol $\mathrm{S}$}}}}. For a string $\bm{w}$}} of length n{{n}}, the algorithm determines whether ${{\bm{w}}}\in{\mathbb{L}}({{\mathcal{G}}})$. To do this, it manipulates items—tuples of the form ${{\mathrm{A}}},{{i}},{{j}}]$, where ${{\mathrm{A}}}\in{{\mathcal{N}}}$ and ${{i}},{{j}}\in{{\left[{{n}}\right]}}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{{\{1,2,\ldots,{{n}}\}}}$. The item ${{\mathrm{A}}},{{i}},{{j}}]$ is realizable if and only if ${{\mathrm{A}}}{{\overset{*}{\Rightarrow}}}{\texttt{w}}_{{i}}{\texttt{w}}_{{{i}}+1}\ldots{\texttt{w}}_{{j}}$, i.e., if there is a sequence of rules that can be applied to the non-terminal $\mathrm{A}$}} that yields ${\texttt{w}}_{{i}}{\texttt{w}}_{{{i}}+1}\ldots{\texttt{w}}_{{j}}$.
@@
Proof.
The proof follows from our definitions. In the base case, if j=i{{j}}={{i}}, then $\mathrm{X}$}} needs to derive exactly the symbol ${\texttt{w}}_{{i}}$ in one step without producing non-terminals (assuming a CFG with no useless non-terminals). In the recursive case, if ${{\mathrm{X}}},{{i}},{{j}}]$ is realizable then there exists some associated parse tree where ${{\mathrm{X}}}{{\overset{*}{\Rightarrow}}}{\texttt{w}}_{{i}}\ldots{\texttt{w}}_{{j}}$. Such a tree can be split by selecting a split node which induces recursive subproblems. If the chosen split node is the root $\mathrm{X}$}}, there exists a rule ${{\mathrm{X}}}\rightarrow{{\mathrm{Y}}}{{\mathrm{Z}}}$}} such that $\mathrm{Y}$}} and $\mathrm{Z}$}} derive disjoint, consecutive substrings of $\bm{w}$}}. If the chosen split node is a non-root ${{\mathrm{Y}}}\in{{\mathcal{N}}}$, then $\mathrm{Y}$}} derives some substring ${\texttt{w}}_{{k}}\ldots{\texttt{w}}_{{l}}$, and $\mathrm{X}$}} derives $\bm{w}$}} where ${\texttt{w}}_{{k}}\ldots{\texttt{w}}_{{l}}$ has been replaced by $\mathrm{Y}$}}. ∎
@@
The proof follows the same structure as the proof of Lem.˜3.1. In the base case, $\mathrm{X}$}} needs to derive in one step the non-terminal $\mathrm{Y}$}} and some non-terminal $\mathrm{Z}$}} such that $\mathrm{Z}$}} derives in one step a symbol at the boundary of ${\texttt{w}}_{{i}}\ldots{\texttt{w}}_{{j}}$ (either ${\texttt{w}}_{{i}}$ or ${\texttt{w}}_{{j}}$). In the recursive case, if ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$ is realizable then there exists a parse tree associated with it where ${{\mathrm{X}}}{{\overset{*}{\Rightarrow}}}{\texttt{w}}_{{i}}\ldots{\texttt{w}}_{{{k}}-1}{{\mathrm{Y}}}{\texttt{w}}_{{{l}}+1}\ldots{\texttt{w}}_{{j}}$. Such a tree can be split by selecting a split node which induces recursive subproblems. If the chosen split node is the root $\mathrm{X}$}}, there exists a rule ${{\mathrm{X}}}\rightarrow{{\mathrm{A}}}{{\mathrm{B}}}$}} such that $\mathrm{A}$}} derives some sentential form ${\texttt{w}}_{{i}}\ldots{\texttt{w}}_{{{k}}-1}{{\mathrm{Y}}}{\texttt{w}}_{{{l}}+1}\ldots p$ and $\mathrm{B}$}} derives the string ${\texttt{w}}_{{p}+1}\ldots{\texttt{w}}_{{j}}$ for some index $p\in{{\left[{{n}}\right]}}$. If the chosen split node is a non-root ${{\mathrm{Z}}}\in{{\mathcal{N}}}$, then $\mathrm{Z}$}} derives the sentential form ${\texttt{w}}_{p}\ldots{\texttt{w}}_{{{k}}-1}{{\mathrm{Y}}}{\texttt{w}}_{{{l}}+1}\ldots{\texttt{w}}_{q}$ and $\mathrm{X}$}} derives the sentential form ${\texttt{w}}_{{{i}}}\ldots{\texttt{w}}_{p-1}{{\mathrm{Z}}}{\texttt{w}}_{q+1}\ldots{\texttt{w}}_{{{j}}}$ for some indices $p,q\in{{\left[{{n}}\right]}}$. ∎
@@
Lemmata˜3.1 and 3.2 state that an item is realizable if it can be decomposed into realizable subproblems. Rather than enumerating all the possible decompositions sequentially, we will leverage parallelism to simultaneously compute the realizability of all the induced subproblems. The term guessing has been coined (chandra-alternation) to denote the ability of a parallel model of computation to attend to a valid computation path given an unbounded set of possible computations via existential branching. By analogy, we can guess which of the correct decompositions of an item is correct by leveraging parallelism, and then recursively verify the induced subproblems in parallel. This suggests natural parallel algorithms for checking realizability, which we present in Algs.˜1 and 2.
@@
1.def $\textsc{Solve}([{{\mathrm{X}}},{{i}},{{j}}])$:
2.if i=j{{i}}={{j}} :
3.  return ${{{{\mathrm{X}}}\rightarrow{\texttt{w}}_{{i}}}}\in{{\mathcal{P}}}$
4.guess an integer $x\in{{\{1,2\}}}$
5.if $x=1$ :
6.  guess a rule ${{{{\mathrm{X}}}\rightarrow{{\mathrm{Y}}}{{\mathrm{Z}}}}}\in{{\mathcal{P}}}$ and ${{k}}\in{{\left[{{n}}\right]}}$
7.  return $\textsc{Solve}([{{\mathrm{Y}}},{{i}},{{k}}-1])\land\textsc{Solve}([{{\mathrm{Z}}},{{k}},{{j}}])$
8.else
9.  guess an item ${{\mathrm{Y}}},{{k}},{{l}}]$
10.  return $\textsc{Solve}([{{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}])$ $\land$ $\textsc{Solve}([{{\mathrm{Y}}},{{k}},{{l}}])$
@@
1.def $\textsc{Solve}([{{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}])$:
2.if ${{k}}={{i}}\land{{l}}={{j}}-1$ :
3.  return $\exists$ ${{\mathrm{Y}}},{{\mathrm{Z}}}\in{{\mathcal{N}}}$ such that ${{{{\mathrm{X}}}\rightarrow{{\mathrm{Y}}}{{\mathrm{Z}}}}}\in{{\mathcal{P}}}$ $\land$ ${{{{\mathrm{Z}}}\rightarrow{\texttt{w}}_{{j}}}}\in{{\mathcal{P}}}$
4.guess an integer $x\in{{\{1,2\}}}$
5.if $x=1$ :
6.  guess a rule ${{{{\mathrm{X}}}\rightarrow{{\mathrm{A}}}{{\mathrm{B}}}}}\in{{\mathcal{P}}}$ and $p\in{{\left[{{n}}\right]}}$
7.  return $\textsc{Solve}([{{\mathrm{A}}},{{i}},p-1]/[{{\mathrm{Y}}},{{k}},{{l}}])\land\textsc{Solve}([{{\mathrm{B}}},p,{{j}}])$
8.else
9.  guess an item ${{\mathrm{Z}}},p,q]$
10.  return $\textsc{Solve}([{{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Z}}},p,q])$ $\land$ $\textsc{Solve}([{{\mathrm{Z}}},p,q]/[{{\mathrm{Y}}},{{k}},{{l}}])$
@@
By definition, ${{\bm{w}}}\in{\mathbb{L}}({{\mathcal{G}}})$ if and only if ${{{{\mathrm{S}}}}},1,{{n}}]$ is realizable. By Lemmata˜3.1 and 3.2, the item ${{{{\mathrm{S}}}}},1,{{n}}]$ is realizable if and only if there exists a decomposition of ${{{{\mathrm{S}}}}},1,{{n}}]$ that respects Lemmata˜3.1 and 3.2. Solve recursively guesses such decompositions, guaranteeing that we will compute a valid decomposition if it exists. ∎
@@
By Thm.˜3.3, for any realizable item, there exists a balanced decomposition of the corresponding parse tree into two trees of roughly equal size which can be represented by two items (the split is at the root) or a slashed item and an item (the split is not at the root). Assuming we can process all possible tree decompositions in parallel, we will necessarily guess the balanced one where subtrees have at most 2n/2+12{{n}}/2+1 nodes (a full binary tree with n{{n}} leaves does not have more than 2n{{n}} nodes). After i{{i}} recursive steps, the current subtrees have at most $\frac{{{n}}}{2^{{{i}}-1}}+{{\mathcal{O}}}(1)$ nodes. Therefore, we will solve all base cases after at most $\log({{n}})+{{\mathcal{O}}}(1)$ steps. ∎
Space complexity.
The bottleneck resides in solving an item ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$, which occupies ${{\mathcal{O}}}({{n}}^{4})$ space, and guessing an item ${{\mathrm{Z}}},p,q]$ that could decompose this problem, which itself occupies ${{\mathcal{O}}}({{n}}^{2})$ space, leading to a total space complexity of ${{\mathcal{O}}}({{n}}^{6})$.
Combining both insights on time- and space-complexity, we can prove the following theorem: See 3.1

### v2

@@ Abstract
Transformers excel empirically on tasks that process well-formed inputs according to some grammar, such as natural language and code. However, it remains unclear how they can process grammatical syntax. In fact, under standard complexity conjectures, standard transformers cannot recognize context-free languages (CFLs), a canonical formalism to describe syntax, or even regular languages, a subclass of CFLs. Past work has shown that ${{\mathcal{O}}}(\log({{n}}))$ looping layers (w.r.t. input length n{{n}}) allows transformers to recognize regular languages, but the question of context-free recognition with looped-transformers remained open. In this work, we show that looped transformers with ${{\mathcal{O}}}(\log({{n}}))$ looping layers and ${{\mathcal{O}}}({{n}}^{6})$ padding tokens can recognize all CFLs. However, training and inference with ${{\mathcal{O}}}({{n}}^{6})$ padding tokens is potentially impractical. Fortunately, we show that, for natural subclasses such as unambiguous CFLs, the recognition problem on transformers becomes more tractable, requiring ${{\mathcal{O}}}({{n}}^{3})$ padding. We empirically validate our results and show that looping helps on a grammar that provably requires logarithmic depth. Overall, our results shed light on the intricacy of CFL recognition by transformers: While general recognition may require an intractable amount of padding, natural constraints such as unambiguity yield efficient recognition algorithms.
@@ 1 Introduction
Transformers are proficient at many natural language (qin2024largelanguagemodelsmeet) and coding (jiang2024surveylargelanguagemodels) tasks, both of which involve processing hierarchical structures. Classically, the ability to process hierarchically nested structures is closely connected to the ability to model context-free languages (CFLs). Analysis of internal representations—syntactic probing—has shown that transformers learn to encode syntactic features relevant for parsing, the task of extracting the syntactic structure of a sentence (hewitt-manning-2019-structural; arps2022probingconstituencystructureneural; zhao-etal-2023-transformers). Moreover, empirical work has found that transformers can leverage parallelism to parse distinct substrings of an input in parallel (allen-zhu2025physics; schulz2025unravelingsyntaxlanguagemodels), suggesting an inherent bias towards learning parallel parsing. However, we lack a precise theory of how CFL recognition can be exactly implemented in model internals, and which types of syntax can transformers provably represent. To this end, we study whether transformers can correctly determine the grammaticality of a sentence according to a context-free grammar.
@@
The problem of determining whether an input is grammatical can be stated as the recognition problem for context-free grammars (CFGs): Given a CFG $\mathcal{G}$}}, can a string $\bm{w}$}} be generated by $\mathcal{G}$}}? Several foundational serial parsing algorithms (earley-parsing; cocke; kasami1965AnER; younger) solve this problem. However, such serial procedures cannot be naturally implemented due to transformers’ highly parallel, fixed-depth structure. Even regular languages, a strict subset of CFLs, cannot be recognized by fixed-depth transformers under the standard complexity conjecture ${\textsc{TC}\textsuperscript{0}}\subsetneq{\textsc{NC}\textsuperscript{1}}$ since regular language recognition is complete for NC1 (barrington_monoids) while fixed-depth transformers fall in TC0 (merrill-etal-2022-saturated; chiang2025transformersuniformtc0). Looping layers help: $\log({{n}})$ looping layers (where n{{n}} is the input length) allow transformers to recognize regular languages (merrill-2025littledepthgoeslong). However, the question of whether logarithmic looping enables CFL recognition remains. In this work, we address it by analyzing the difficulty of recognizing various CFL classes by transformers. We conceptualize the difficulty in terms of extra resources needed: looping layers and appending blank padding tokens (merrill2025exactexpressivepowertransformers).
@@
While general CFL recognition cannot be implemented by fixed-depth transformers under standard complexity conjectures, our first result shows via a direct construction that it can be expressed by looping layers ${{\mathcal{O}}}(\log({{n}}))$ times and with ${{\mathcal{O}}}({{n}}^{6})$ padding tokens. To the best of our knowledge, this constitutes the first proof of general CFL recognition by transformers. We then ask whether simpler classes of CFLs can be recognized by transformers with fewer resources. We find that the answer is affirmative: We show that natural subclasses of CFLs can be recognized by simpler transformers. In particular, we identify unambiguity and linearity as two natural properties that make CFL recognition more tractable. Unambiguous CFLs, characterized by strings having at most one possible parse, allow for recognition with reduced padding but more looping. This aligns with transformers’ struggles to parse ambiguous grammars in practice (khalighinejad-etal-2023-approximating). Furthermore, imposing linearity (where each grammar rule has at most one non-terminal on its right-hand side) reduces the amount of looping and padding required for recognizing unambiguous CFLs. We empirically test when looping helps generalization and find it to increase the performance on a CFL known to require ${{\mathcal{O}}}(\log({{n}}))$ time on parallel models of computation, namely the language of variable-free Boolean formulas (buss-bfvp).
@@
In summary, we leverage theory on parallel recognition of CFLs to show that logarithmically-looped transformers can recognize CFLs, characterizing the padding requirements for different relevant subclasses. The results imply that, in contrast to the depth expected to implement a serial algorithm such as CKY, exponentially less depth suffices to recognize general CFLs. While this comes with increased space (padding) requirements in the general case, we show the space can be reduced for natural CFL subclasses. Our main results are summarized in Tab.˜1.
@@ 2.1 Context-free Grammars
The language of a CFG $\mathcal{G}$}} is the set ${\mathbb{L}}({{\mathcal{G}}})\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}\{{{\bm{w}}}\in{{{{\Sigma}}^{*}}}\mid{{{{\mathrm{S}}}}}{{\overset{*}{\Rightarrow}}}{{\bm{w}}}\}$.
@@ 2.2 Transformers
We consider the idealization of transformers from merrill-2025littledepthgoeslong; merrill2025exactexpressivepowertransformers. In short, [^1] We refer to App. A for more details on the transformer model. we study average hard attention transformers (AHATs), where the attention normalization function returns a uniform average of the values of tokens that maximize the attention score. In practice, training dynamics lead attention heads in transformers to approximate average-hard attention (merrill-etal-2021-effects). The transformers use multi-pre-norm, where the layer normalization is applied before the residual connection on either the entire hidden state or on distinct subsets thereof (merrill2024expressivepowertransformerschain). We further assume logarithmic-precision arithmetic, where computations are performed with ${{\mathcal{O}}}(\log({{n}}))$ bits for an input of size n{{n}}. Coupling AHATs and log-precision unlocks useful gadgets such as storing string indices, counting symbol occurrences across the string and performing equality checks of values stored in residual streams at separate positions (merrill2024expressivepowertransformerschain). We assume input strings to the transformer are augmented with both a beginning-of-sequence (bos) and end-of-sequence (eos) token. Denote by ${{{\bm{x}}}}_{{{\textsc{eos}}}}^{L}$ the contextual representation of eos at end of the forward pass of the transformer. We apply a linear classifier to ${{{\bm{x}}}}_{{{\textsc{eos}}}}^{L}$ to determine string acceptance.
@@ 3 Recognizing General CFLs with Transformers
Our goal is to recognize a CFL represented by a grammar in CNF (Def.˜2.4) with start symbol $\mathrm{S}$}}}}. For a string $\bm{w}$}} of length n{{n}}, the algorithm determines whether ${{\bm{w}}}\in{\mathbb{L}}({{\mathcal{G}}})$. To do this, it manipulates items—tuples of the form ${{\mathrm{A}}},{{i}},{{j}}]$, where ${{\mathrm{A}}}\in{{\mathcal{N}}}$ and ${{i}},{{j}}\in{{\left[{{n}}\right]}}\mathrel{\stackrel{{\scriptstyle\textnormal{\tiny def}}}{{=}}}{{\{1,2,\ldots,{{n}}\}}}$. The item ${{\mathrm{A}}},{{i}},{{j}}]$ is realizable if and only if ${{\mathrm{A}}}{{\overset{*}{\Rightarrow}}}{\texttt{w}}_{{i}}{\texttt{w}}_{{{i}}+1}\ldots{\texttt{w}}_{{j}}$, i.e., if there is a sequence of rules that can be applied to the non-terminal $\mathrm{A}$}} that yields ${\texttt{w}}_{{i}}{\texttt{w}}_{{{i}}+1}\ldots{\texttt{w}}_{{j}}$.
@@
We can also devise a recursive procedure to evaluate the realizability of a slashed items.
@@
Lem.˜3.1 and Lem.˜3.2 both follow from first principles. In the base case of ${{\mathrm{X}}},{{i}},{{j}}]$, if j=i{{j}}={{i}}, then $\mathrm{X}$}} needs to derive exactly the symbol ${\texttt{w}}_{{i}}$ in one step without producing non-terminals (assuming a CFG with no useless non-terminals) because the grammar is in CNF. Similarly, in the base case of ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$, $\mathrm{X}$}} needs to derive in one step a non-terminal $\mathrm{Y}$}} and a non-terminal $\mathrm{Z}$}} such that $\mathrm{Z}$}} derives in one step a symbol at the boundary of ${\texttt{w}}_{{i}}\ldots{\texttt{w}}_{{j}}$. In the recursive case, a (slashed) item exists if and only if it can be described by a parse tree. Such a parse tree can be broken down into recursive subproblems by selecting a split node, which induces two sub-items to solve. We distinguish two equivalent cases: selecting the root as the split (which requires replacing the root $\mathrm{X}$}} with non-terminals it can produce in one-step) [^2] This is the same recursive rule as allen-zhu2025physics’s recursive formula., or a non-root node as the split. ∎
@@
Lemmata˜3.1 and 3.2 state that an item is realizable if it can be decomposed into realizable subproblems. Rather than enumerating all the possible decompositions sequentially, we will leverage parallelism to simultaneously compute the realizability of all the induced subproblems. The term guessing has been coined (chandra-alternation) to denote the ability of a parallel model of computation to attend to a valid computation path given an unbounded set of possible computations via parallel branching. By analogy, we can guess which of the correct decompositions of an item is correct by leveraging parallelism, and then recursively verify the induced subproblems in parallel. This suggests natural parallel algorithms for checking realizability, which we present in Algs.˜1 and 2.
@@
[…]
@@
[…]
@@
By definition, ${{\bm{w}}}\in{\mathbb{L}}({{\mathcal{G}}})$ if and only if ${{{{\mathrm{S}}}}},1,{{n}}]$ is realizable. By Lemmata˜3.1 and 3.2, the item ${{{{\mathrm{S}}}}},1,{{n}}]$ is realizable if and only if there exists a decomposition of ${{{{\mathrm{S}}}}},1,{{n}}]$ that respects Lemmata˜3.1 and 3.2. Solve recursively computes all such decompositions, guaranteeing that we will encounter a valid decomposition if it exists. ∎
@@
By Thm.˜3.3, for any realizable item, there exists a balanced decomposition of the corresponding parse tree into two trees of roughly equal size which can be represented by two items (the split is at the root) or a slashed item and an item (the split is not at the root). Assuming we can process all possible tree decompositions in parallel, we will necessarily guess the balanced one where subtrees have at most 2n/2+12{{n}}/2+1 nodes (a full binary tree with n{{n}} leaves does not have more than 2n{{n}} nodes). After i{{i}} recursive steps, the current subtrees have at most $\frac{{{n}}}{2^{{{i}}-1}}+{{\mathcal{O}}}(1)$ nodes. Therefore, we will solve all items after at most $\log({{n}})+{{\mathcal{O}}}(1)$ steps. ∎
Space complexity. The bottleneck resides in solving an item ${{\mathrm{X}}},{{i}},{{j}}]/[{{\mathrm{Y}}},{{k}},{{l}}]$, which occupies ${{\mathcal{O}}}({{n}}^{4})$ space, and guessing an item ${{\mathrm{Z}}},p,q]$ that could decompose this problem, which itself occupies ${{\mathcal{O}}}({{n}}^{2})$ space, leading to a total space complexity of ${{\mathcal{O}}}({{n}}^{6})$. Combining both insights on time- and space-complexity, we can then prove the following theorem: See 3.1